[Brahms-l] Shift report 20050220 00:00-08:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Sun Feb 20 2005 - 08:46:22 EST
Shifter leader was SZ
Shifters were 


1. The Machine worked properly.
2. RUNS:

| RunNo | tr2     | tr3     |
+-------+---------+---------+
| 13310 | 213.56K | 162.93K |
| 13311 | 131.51K | 136.88K |
| 13312 | 19.36K  | 24.87K  |
| 13313 | 252.97K | 214.70K |
----------------   ---------|
| TOTAL | 828.15K | 660.98K |
+--------+--------+----------
3. T1 have some problem. 
At the and of shift Hiro is working to fixed!

E-log:
http://brahms-web.brahms.bnl.gov/cgi-bin/elog.pl?nb=2005&action=view&page=128&scroll=true


_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sun Feb 20 08:46:52 2005

This archive was generated by hypermail 2.1.8 : Sun Feb 20 2005 - 08:47:05 EST